3. The owner of a cell phone store at Greenhill Shopping Center was not happy when he learned that the

average number of cell phones sold daily was only 8 mobile phones. He hired a new seller and told him to

tally the daily sales for 20 days. The following are the number of cell phones sold for the past 20 days.

9

11

12

6

12

8

17

14

10

12

11

18

15

14

17

Does the collected data present sufficient evidence to indicate that the average number of cell phones sold

daily has increased? Use 0.05 level of significance. Assume normality in the population.


1
Expert's answer
2021-07-16T13:41:09-0400
"9, 9, 9, 8, 9, 9, 11, 12, 12, 16,""10, 12, 8, 11, 17, 18, 15, 14, 17,14""n=20, \sum_ix_i=240""\bar{x}=\dfrac{1}{20}\sum_ix_i=\dfrac{240}{20}=12""\sum_i(x_i-\bar{x})^2=(9-12)^2+(9-12)^2+(9-12)^2""+(8-12)^2+(9-12)^2+(9-12)^2+(11-12)^2""+(12-12)^2+(12-12)^2+(16-12)^2+(10-12)^2""+(12-12)^2+(8-12)^2+(11-12)^2+(17-12)^2""+(18-12)^2+(15-12)^2+(14-12)^2+(17-12)^2""+(14-12)^2=202""s^2=\dfrac{\sum_i(x_i-\bar{x})^2}{n-1}=\dfrac{202}{20-1}=\dfrac{202}{19}""\approx10.631579""s=\sqrt{s^2}\approx3.260610"

Hypothesized Population Mean "\mu=8"

Sample Standard Deviation "s=3.260610"

Sample Size "n=20"

Sample Mean "\bar{x}=12"

Significance Level "\alpha=0.05"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu\leq8"

"H_1: \mu>8"


This corresponds to right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


Based on the information provided, the significance level is "\alpha=0.05," "df=n-1=19" degrees of freedom and the critical value for right-tailed test is "t_c=1.729133."


The rejection region for this right-tailed test is "R=\{t:t>1.729133\}."


The "t" - statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{12-8}{3.260610/\sqrt{20}}\approx5.486257"

Since it is observed that "t=5.486257>1.729133=t_c," it is then concluded that the null hypothesis is rejected.


Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "8," at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value for right-tailed, the significance level 

"\alpha=0.05, t=5.486257, df=19" is "p=0.000014," and since "p=0.000014<0.05=\alpha," it is concluded that the null hypothesis is rejected.


Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "8," at the "\alpha=0.05" significance level.